OpenStudy (anonymous):

Find values of s: \[\frac{s+\sqrt{s^2+2}}{2}\leq1\]

OpenStudy (amistre64):

the radical has no restrictions ...

OpenStudy (anonymous):

wat do u mean? Group Master?

OpenStudy (amistre64):

if we do this normally ..... as if theres another way s+sqrt(s^2 +2) </ 2

OpenStudy (amistre64):

s aint got restrictions under the radical since when you square it it goes +

OpenStudy (amistre64):

sqrt(s^2 +2) </ 2-s

OpenStudy (anonymous):

answer would be s>=1/2 amistre

OpenStudy (anonymous):

sorry s<=1/2

OpenStudy (amistre64):

s^2 +2 </ 4 +s^2 -4s 2 </ 4-4s -2 </ -4s 2>/ 4s 1/2 >/ s

OpenStudy (amistre64):

yep; thats what I get :) without a dbl chk

OpenStudy (anonymous):

\[\sqrt{s^2+2}\leq2-s \implies s^2+2\leq(s-2)^2\ for\ positive\ s.\]

OpenStudy (anonymous):

\(\frac{-7}{4}\le s \le \frac{1}{2}\). That's what I got.

OpenStudy (anonymous):

For negative s, let x=-s and \[\sqrt{x^2+2}\leq2+x\implies x^2+2\leq4+x^2+4x\implies-2\leq4x\implies x\geq-1/2\implies s\leq1/2\]

OpenStudy (anonymous):

i got x < 1/2

OpenStudy (anonymous):

u get s<=1/2 always.

OpenStudy (anonymous):

oops so did everyone else!

OpenStudy (anonymous):

\[(\sqrt{s ^{2}+2})^{2}\le (2-s)^{2}\] \[s^{2}+2\le4-4s+s ^{2}\] \[4s-2\le0\] \[s \le1/2 \forall s \in \mathbb{R}\]

OpenStudy (anonymous):

\(s\le\frac{1}{2}\) is not enough condition. For example s=0 gives that \(\frac{\sqrt{2}}{2}\le\frac{1}{2}\), which is not right. So, I think that \(s\le-\frac{7}{4}\).